OpenStudy (anonymous):

Subtract: (7x+3)-(2x+1)

OpenStudy (anonymous):

5x+2

OpenStudy (uri):

@numinumi Just remove the parenthesis.

OpenStudy (anonymous):

7x+3-2x-1 =(7-2)x+3-1 =5x+2

OpenStudy (uri):

Always remember, double negative =positive. negative + positive = negative.